National Defence Academy (NDA) · Topic Collection

Quantitative Aptitude PYQs

Arithmetic, number systems, percentages, ratios, profit and loss, time and work, algebra, geometry, mensuration, and trigonometry.

2520 Total Questions21 Official PapersYears: 2015–2025Search in Quantitative Aptitude
Showing questions 1421–1440 of 2520Page 72 of 100
NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1421

National Defence Academy (NDA) 2019• Mathematics• Mathematics

From a deck of cards, cards are taken out with replacement. What is the probability that the fourteenth card taken out is an ace?

  1. AOption (a)

    \(\frac{1}{{51}}\)

  2. BOption (b)

    \(\frac{4}{{51}}\)

  3. COption (c)

    \(\frac{1}{{52}}\)

  4. DOption (d)

    \(\frac{1}{{13}}\)

View answer & explanation

Correct answer

Answer: (d)D.

\(\frac{1}{{13}}\)

Why this answer is correct

Calculation

Total number of arrangements of 52 cards is 52!.

Now, Lets consider the 14th position.

The 14th position can be filled with ace in 4 ways.

After the 14th position is filled, remaining 51 positions can be filled with the remaining 51 cards in 51! ways.

So total favourable arrangements is 4 × 51!

Required probability \(= {\rm{}}\frac{{4{\rm{;}} \times {\rm{;}}51!}}{{52!}} = \frac{1}{{13}}\)

Correct option is (4).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1422

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If A and B are two events such that P(A) = 0.5, P(B) = 0.6 and P(A ∩ B) = 0.4, then what is \(P\left( {\overline {A \cup B} } \right)\) equal to?

  1. AOption (a)

    0.9

  2. BOption (b)

    0.7

  3. COption (c)

    0.5

  4. DOption (d)

    0.3

View answer & explanation

Correct answer

Answer: (d)D.

0.3

Why this answer is correct

Concept

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

\({\rm{P}}\left( {{\rm{A}} \cup {\rm{B}}} \right) + {\rm{P}}\left( {\overline {{\rm{A}} \cup {\rm{B}}} } \right) = 1\)

Calculation

P(A) = 0.5, P(B) = 0.6, P(A ∩ B) = 0.4

⇒ P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ P(A ∪ B) = 0.5 + 0.6 – 0.4

⇒ P(A ∪ B) = 0.7

\(⇒ {\rm{P}}\left( {\overline {{\rm{A}} \cup {\rm{B}}} } \right) = 1 - {\rm{P}}\left( {\ {{\rm{A}} \cup {\rm{B}}} } \right)\)

⇒ 1 - P( A ∪ B) = 1 – 0.7

\(⇒ {\rm{P}}\left( {\overline {{\rm{A}} \cup {\rm{B}}} } \right) = 0.3\)

Correct option is (4).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1423

National Defence Academy (NDA) 2019• Mathematics• Mathematics

A problem is given to three students A, B and C whose probabilities of solving the problem are \(\frac{1}{2},\frac{3}{4}\) and \(\frac{1}{4}\) respectively. What is the probability that the problem will be solved if they all solve the problem independently?

  1. AOption (a)

    \(\frac{{29}}{{32}}\)

  2. BOption (b)

    \(\frac{{27}}{{32}}\)

  3. COption (c)

    \(\frac{{25}}{{32}}\)

  4. DOption (d)

    \(\frac{{23}}{{32}}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{{29}}{{32}}\)

Why this answer is correct

Calculation

\({\rm{P}}\left( {\rm{A}} \right) = \frac{1}{2},{\rm{;;P}}\left( {\rm{B}} \right) = \frac{3}{4},{\rm{;;P}}\left( {\rm{C}} \right) = \frac{1}{4}{\rm{;}}\)

 All students solves the problem independently, so if at least one of them solves the problem then the problem is considered solved.

Probability that the problem is unsolved

= [1 – P(A)][1 – P(B)][1 – P(C)]

= \(\frac{3}{{32}}\)

The probability that the problem will be solved

= 1 – Probability that the problem is unsolved

\( = \frac{{29}}{{32}}\)

Correct option is (1).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1424

National Defence Academy (NDA) 2019• Mathematics• Mathematics

A pair of fair dice is rolled. What is the probability that the second dice lands on a higher value than does the first?

  1. AOption (a)

    \(\frac{1}{4}\)

  2. BOption (b)

    \(\frac{1}{6}\)

  3. COption (c)

    \(\frac{5}{{12}}\)

  4. DOption (d)

    \(\frac{5}{{18}}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{5}{{12}}\)

Why this answer is correct

Concept

The sample space when a pair of dice is rolled is as follows:

(1, 1), (1, 2), (1, 3), (1, 4), (1, 5), (1, 6)

(2, 1), (2, 2), (2, 3), (2, 4), (2, 5), (2, 6)

(3, 1), (3, 2), (3, 3), (3, 4), (3, 5), (3, 6)

(4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6)

(5, 1), (5, 2), (5, 3), (5, 4), (5, 5), (5, 6)

(6, 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6)

Total outcomes = 36

Calculation

The pairs in which the second number is greater than the first number are as follows:

(1, 2), (1, 3), (1, 4), (1, 5), (1, 6), (2, 3), (2, 4), (2, 5), (2, 6), (3, 4), (3, 5), (3, 6), (4, 5), (4, 6), (5, 6)

Total favourable outcomes = 15

Required probability

\( = \frac{{15}}{{36}} = \frac{5}{{12}}\)

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1425

National Defence Academy (NDA) 2019• Mathematics• Mathematics

A fair coin is tossed and an unbiased dice is rolled together. What is the probability of getting a 2 or 4 or 6 along with head?

  1. AOption (a)

    \(\frac{1}{2}\)

  2. BOption (b)

    \(\frac{1}{3}\)

  3. COption (c)

    \(\frac{1}{4}\)

  4. DOption (d)

    \(\frac{1}{6}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{1}{4}\)

Why this answer is correct

Concept

The sample space when a fair coin is tossed and an unbiased dice is rolled together is as follows:

(H, 1), (H, 2), (H, 3), (H, 4), (H, 5), (H, 6)

(T, 1), (T, 2), (T, 3), (T, 4), (T, 5), (T, 6)

Total outcomes = 12

Calculation

Pairs in which we get a 2 or 4 or 6 along with head are as follows:

(H, 2), (H, 4), (H, 6)

Total favourable outcomes = 3

Required probability

\(= \frac{3}{{12}}\)

\(= \frac{1}{{4}}\)

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1426

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If A, B, C are three events, then what is the probability that at least two of these events occur together?

  1. AOption (a)

    P(A ∩ B) + P(B ∩ C) + P(C ∩ A)

  2. BOption (b)

    P(A ∩ B) + P(B ∩ C) + P(C ∩ A) – P(A ∩ B ∩ C)

  3. COption (c)

    P(A ∩ B) + P(B ∩ C) + P(C ∩ A) – 2P(A ∩ B ∩ C)

  4. DOption (d)

    P(A ∩ B) + P(B ∩ C) + P(C ∩ A) – 3P(A ∩ B ∩ C)

View answer & explanation

Correct answer

Answer: (c)C.

P(A ∩ B) + P(B ∩ C) + P(C ∩ A) – 2P(A ∩ B ∩ C)

Why this answer is correct

Concept

Calculation

Consider the following Venn diagram:

Atleast two of the events means exactly two events will occur or all the three events will occur.

Region b, d and f represents exactly two events.

b = (A ∩ B) - (A ∩ B ∩ C)

d = (A ∩ C) - (A ∩ B ∩ C)

f = (B ∩ C) - (A ∩ B ∩ C)

Region e represents all the three events.

e = (A ∩ B ∩ C)

The probability that at least two of these events occur together

= P(b) + P(d) + P(f) + P(e)

= P(A ∩ B) + P(A ∩ C) + P(B ∩ C) – 2P(A ∩ B ∩ C)

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1427

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If two variables X and Y are independent, then what is the correlation coefficient between them?

  1. AOption (a)

    1

  2. BOption (b)

    -1

  3. COption (c)

    0

  4. DOption (d)

    None of the above

View answer & explanation

Correct answer

Answer: (c)C.

0

Why this answer is correct

Concept

If X and Y are independent, then E(XY) = E(X)E(Y)

Covariance Cov(X, Y) of X and Y = E(XY) – E(X)E(Y)

\(Correlation~coefficient = \frac{{{\bf{Cov}}\left( {{\bf{X}},;{\bf{Y}}} \right)}}{{{{\bf{\sigma }}_{\bf{X}}}{{\bf{\sigma }}_{\bf{Y}}}}}{\rm{;}}\)

Calculation

Given that X and Y are independent, then the Covariance Cov(X, Y) = 0

Hence, the correlation coefficient = 0.

The correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1428

National Defence Academy (NDA) 2019• Mathematics• Mathematics

Two independent events A and B are such that \({\rm{P}}\left( {{\rm{A}} \cup {\rm{B}}} \right) = \frac{2}{3}\) and \({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{6}\). If P(B) < P(A), then what is P(B) equal to?

  1. AOption (a)

    \(\frac{1}{4}\)

  2. BOption (b)

    \(\frac{1}{3}\)

  3. COption (c)

    \(\frac{1}{2}\)

  4. DOption (d)

    \(\frac{1}{6}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{1}{3}\)

Why this answer is correct

Concept

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

If A and B are independent, then P(A ∩ B) = P(A)P(B).

Calculation

\({\rm{P}}\left( {{\rm{A}} \cup {\rm{B}}} \right) = \frac{2}{3},{\rm{;P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{6}\)

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ P(A) + P(B) = P(A ∪ B) + P(A ∩ B)

\( \Rightarrow {\rm{P}}\left( {\rm{A}} \right) + {\rm{;P}}\left( {\rm{B}} \right) = \frac{5}{6}\)

Given A and B are independent so, P(A ∩ B) = P(A)P(B)

\( \Rightarrow {\rm{P}}\left( {\rm{A}} \right) = \frac{1}{{6{\rm{P}}\left( {\rm{B}} \right)}}\)

As, \({\rm{P}}\left( {\rm{A}} \right) + {\rm{;P}}\left( {\rm{B}} \right) = \frac{5}{6}\)

\( \Rightarrow ;\frac{1}{{6{\rm{P}}\left( {\rm{B}} \right)}} + {\rm{;P}}\left( {\rm{B}} \right) = \frac{5}{6}\)

⇒ 6P(B)2 - 5P(B) + 1 = 0

\(\Rightarrow {\rm{P}}\left( {\rm{B}} \right) = \frac{{5 \pm \sqrt {{{\left( { - 5} \right)}^2} - 4 \times 6} }}{{2 \times 6}}\)

\(\Rightarrow {\rm{P}}\left( {\rm{B}} \right) = {\rm{;}}\frac{1}{2},{\rm{;}}\frac{1}{3}\)

\({\rm{P}}\left( {\rm{A}} \right) = \frac{5}{6} - {\rm{P}}\left( {\rm{B}} \right)\)

\(\Rightarrow {\rm{P}}\left( {\rm{A}} \right) = \frac{1}{3},{\rm{;}}\frac{1}{2}\)

Given that P(B) < P(A) so the possible value of P(B) is \(\frac{1}{3}\)

Correct option is (2).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1429

National Defence Academy (NDA) 2019• Mathematics• Mathematics

The mean of 100 observations is 50 and the standard deviation is 10. If 5 is subtracted from each observation and then it is divided by 4, then what will be the new mean and the new standard deviation respectively?

  1. AOption (a)

    45, 5

  2. BOption (b)

    11.25, 1.25

  3. COption (c)

    11.25, 2.5

  4. DOption (d)

    12.5, 2.5

View answer & explanation

Correct answer

Answer: (c)C.

11.25, 2.5

Why this answer is correct

Concept

Mean of n observations = \(\frac{1}{{\bf{n}}}\mathop \sum \limits_{{\bf{i}} = 1}^{{\bf{i}} = {\bf{n}}} {{\bf{X}}_{\bf{i}}}\)

Standard deviation = \(\sqrt {\frac{1}{{\rm{n}}}\mathop \sum \limits_{{\bf{i}} = 1}^{{\bf{i}} = {\bf{n}}} {{\left( {{{\bf{X}}_{\bf{i}}} - {\bf{mean}}} \right)}^2}} \)

where, Xi = observation

Calculation

The mean of 100 observations is 50

\( \Rightarrow \frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = 100} {{\rm{X}}_{\rm{i}}} = 50\)

\( \Rightarrow \mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = 100} {{\rm{X}}_{\rm{i}}} = 5000\)

Standard deviation is 10

\(\Rightarrow \sqrt {\frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\left( {{{\rm{X}}_{\rm{i}}} - 50} \right)}^2}} = 10\)

5 is subtracted from each Xi and the resultant is divided by 4 so, the new observation becomes \(\frac{{{{\bf{X}}_{\bf{i}}} - 5}}{4}\).

Mean of new 100 observations is

\(= \frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = 100} \left( {\frac{{{{\rm{X}}_{\rm{i}}} - 5}}{4}} \right)\)

\(= \frac{1}{{100}}\left[ {\frac{1}{4}\mathop \sum \limits_{i = 1}^{i = 100} {{\rm{X}}_{\rm{i}}} - \frac{1}{4}\mathop \sum \limits_{i = 1}^{i = 100} 5} \right]\)

\(= \frac{1}{{100}}\left[ {\frac{{5000}}{4} - \frac{{100 \times 5}}{4}} \right]\)

= 11.25

Standard deviation of new observations is

\(= \sqrt {\frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\left( {\frac{{{{\rm{X}}_{\rm{i}}} - 5}}{4} - 11.25} \right)}^2}} \)

\(= \sqrt {\frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\left( {\frac{{{{\rm{X}}_{\rm{i}}} - 50}}{4}} \right)}^2}}\)

\(= \sqrt {\frac{1}{{1600}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\left( {{{\rm{X}}_{\rm{i}}} - 50} \right)}^2}}\)

\(= \frac{1}{4}\sqrt {\frac{1}{{100}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\left( {{{\rm{X}}_{\rm{i}}} - 50} \right)}^2}}\)

\(= \frac{{10}}{4}\)

= 2.5

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1430

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If two fair dice are rolled then what is the conditional probability that the first dice lands on 6 given that the sum of numbers on the dice is 8?

  1. AOption (a)

    \(\frac{1}{3}\)

  2. BOption (b)

    \(\frac{1}{4}\)

  3. COption (c)

    \(\frac{1}{5}\)

  4. DOption (d)

    \(\frac{1}{6}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{1}{5}\)

Why this answer is correct

Concept

If A and B are two events, then the conditional probability \({\bf{P}}\left( {{\bf{A}}/{\bf{B}}} \right) = \frac{{{\bf{P}}\left( {{\bf{A}}; \cap ;{\bf{B}}} \right)}}{{{\bf{P}}\left( {\bf{B}} \right)}}\)

The sample space when a pair of dice is rolled is as follows:

(1, 1), (1, 2), (1, 3), (1, 4), (1, 5), (1, 6)

(2, 1), (2, 2), (2, 3), (2, 4), (2, 5), (2, 6)

(3, 1), (3, 2), (3, 3), (3, 4), (3, 5), (3, 6)

(4, 1), (4, 2), (4, 3), (4, 4), (4, 5), (4, 6)

(5, 1), (5, 2), (5, 3), (5, 4), (5, 5), (5, 6)

(6, 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6)

Total outcomes = 36

Calculation

Let A be the set of all pairs in which 6 appears on the first dice and B be the set of all pairs sum of whose numbers is 8.

A = (6, 1), (6, 2), (6, 3), (6, 4), (6, 5), (6, 6)

B = (2, 6), (3, 5), (4, 4), (5, 3), (6, 2)

A ∩ B = (6, 2)

\({\rm{P}}\left( {{\rm{A}} \cap {\rm{B}}} \right) = \frac{1}{{36}}\)

\({\rm{P}}\left( {\rm{B}} \right) = \frac{5}{{36}}\)

Required probability is \({\bf{P}}\left( {{\bf{A}}/{\bf{B}}} \right) = \frac{{{\bf{P}}\left( {{\bf{A}}; \cap ;{\bf{B}}} \right)}}{{{\bf{P}}\left( {\bf{B}} \right)}} = \frac{1}{5}\)

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1431

National Defence Academy (NDA) 2019• Mathematics• Mathematics

Two symmetric dice flipped with each dice having two sides painted red, two painted black, one painted yellow and the other painted white. What is the probability that both land on the same colour?

  1. AOption (a)

    \(\frac{3}{{18}}\)

  2. BOption (b)

    \(\frac{2}{9}\)

  3. COption (c)

    \(\frac{5}{{18}}\)

  4. DOption (d)

    \(\frac{1}{3}\)

View answer & explanation

Correct answer

Answer: (c)C.

\(\frac{5}{{18}}\)

Why this answer is correct

Concept

The sample space is as follows:

(R1, r1), (R1, r2), (R1, b1), (R1, b2), (R1, y), (R1, w)

(R2, r1), (R2, r2), (R2, b1), (R2, b2), (R2, y), (R2, w)

(B1, r1), (B1, r2), (B1, b1), (B1, b2), (B1, y), (B1, w)

(B2, r1), (B2, r2), (B2, b1), (B2, b2), (B2, y), (B2, w)

(Y, r1), (Y, r2), (Y, b1), (Y, b2), (Y, y), (Y, w)

(W, r1), (W, r2), (W, b1), (W, b2), (W, y), (W, w)

R1, R2, B1, B2, Y, W are black, red, yellow and white faces of dice 1.

r1, r2, b1, b2, y, w are black, red, yellow and white faces of dice 2.

Total outcomes = 36

Calculation

Pairs with same colours are as follows:

(R1, r1), (R1, r2), (R2, r1), (R2, r2), (B1, b1), (B1, b2), (B2, b1), (B2, b2), (Y, y), (W, w)

Number of Favourable outcomes = 10

Required Probability = \(\frac{{10}}{{36}} = \frac{5}{{18}}\)

Correct option is (3).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1432

National Defence Academy (NDA) 2019• Mathematics• Mathematics

There are n socks in a drawer, of which 3 socks are red. If 2 of the socks are chosen randomly and the probability that both selected socks are red is  \(\frac{1}{2}\), then what is the value of n?

  1. AOption (a)

    3

  2. BOption (b)

    4

  3. COption (c)

    5

  4. DOption (d)

    6

View answer & explanation

Correct answer

Answer: (b)B.

4

Why this answer is correct

Concept

\({\bf{C}}\left( {{\bf{n}},;{\bf{r}}} \right) = ;\frac{{{\bf{n}}!}}{{{\bf{r}}!\left( {{\bf{n}} - {\bf{r}}} \right)!}}\)

Calculation

From n socks, 2 can be chosen in C(n, 2) ways.

From 3 red socks, 2 can be chosen in C(3, 2) ways

Probability that both socks selected are red

\( = \frac{{{\rm{C}}\left( {3,{\rm{;;;}}2} \right)}}{{{\rm{C}}\left( {{\rm{n}},{\rm{;;;}}2} \right)}}\)

\(= \frac{6}{{{\rm{n}}\left( {{\rm{n}} - 1} \right)}}\)

It is given that Probability that both socks selected are red = \(\frac{1}{2}\)

\( \Rightarrow ;\frac{6}{{{\rm{n}}\left( {{\rm{n}} - 1} \right)}} = \frac{1}{2}\)

⇒ n2 – n – 12 = 0

⇒ (n - 4)(n + 3) = 0

⇒ n = 4, -3

n cannot be negative so, n = 4.

Correct option is (2).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1433

National Defence Academy (NDA) 2019• Mathematics• Mathematics

Two cards are chosen at random from a deck of 52 playing cards. What is the probability that both of them have the same value?

  1. AOption (a)

    \(\frac{1}{{17}}\)

  2. BOption (b)

    \(\frac{3}{{17}}\)

  3. COption (c)

    \(\frac{5}{{17}}\)

  4. DOption (d)

    \(\frac{7}{{17}}\)

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{{17}}\)

Why this answer is correct

Concept

\({\bf{C}}\left( {{\bf{n}},;{\bf{r}}} \right) = \frac{{{\bf{n}}!}}{{{\bf{r}}!\left( {{\bf{n}} - {\bf{r}}} \right)!}}\)

Calculation

From 52 cards, 2 can be chosen in C(52, 2) ways.

There are 4 suits in a deck of 52 cards and each number is present in all the 4 suits.

There are 13 numbers form 1(Ace card) to 13(King card).

If we consider the number 1, then from 4 Aces, 2 can be chosen in C(4, 2) ways.

There are 13 numbers in total so, total ways in which two same numbers can be chosen is 13 × C(4, 2)

Required probability

\( = \frac{{13; \times ;{\rm{C}}\left( {4,{\rm{;}}2} \right)}}{{{\rm{C}}\left( {52,{\rm{;}}2} \right)}}\)

\(= \frac{{13; \times ;12}}{{52; \times ;51}}\)

\(= \frac{1}{{17}}\)

Correct option is (1).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1434

National Defence Academy (NDA) 2019• Mathematics• Mathematics

In eight throws of a die, 5 or 6 is considered a success. The mean and standard deviation of total number of successes is respectively given by

  1. AOption (a)

    \(\frac{8}{3},\frac{{16}}{9}\)

  2. BOption (b)

    \(\frac{8}{3},\frac{4}{3}\)

  3. COption (c)

    \(\frac{4}{3},\frac{4}{3}\)

  4. DOption (d)

    \(\frac{4}{3},\frac{{16}}{9}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{8}{3},\frac{4}{3}\)

Why this answer is correct

Concept

When a fair dice is rolled, probability of getting any number is \(\frac{1}{6}\).

Suppose an experiment is repeated n times. Probability of success be p.

If X denote the number of success, then

P(X = k) = C(n, k) × pk (1 – p)n - k , 0 ≤ k ≤ n

where

\({\bf{C}}\left( {{\bf{n}},;{\bf{k}}} \right) = ;\frac{{{\bf{n}}!}}{{{\bf{k}}!\left( {{\bf{n}} - {\bf{k}}} \right)!}}\)

Mean of X = np

Standard deviation of X \(= \sqrt {{\bf{np}}\left( {1 - {\bf{p}}} \right)}\)

Calculation

A dice is rolled 8 times so, n = 8.

In a single roll of dice, probability of getting 5 or 6 is

= probability of getting 5 + probability of getting 6

\(= \frac{1}{6} + \frac{1}{6}\)

\(= \frac{1}{3}\)

So probability of success \({\bf{p}} = \frac{1}{3}\)

\({\bf{Mean}} = {\bf{n}} \times {\bf{p}} = \frac{8}{3}\)

\({\bf{Standard}};{\bf{deviation}} = \sqrt {{\bf{np}}\left( {1 - {\bf{p}}} \right)} = \frac{4}{3}\)

Correct option is (2).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1435

National Defence Academy (NDA) 2019• Mathematics• Mathematics

A and B are two events such that A̅ and B̅ are mutually exclusive. If P(A) = 0.5 and P(B) = 0.6, then what is the value of P(A|B)?

  1. AOption (a)

    \(\frac{1}{5}\)

  2. BOption (b)

    \(\frac{1}{6}\)

  3. COption (c)

    \(\frac{2}{5}\)

  4. DOption (d)

    \(\frac{1}{3}\)

View answer & explanation

Correct answer

Answer: (b)B.

\(\frac{1}{6}\)

Why this answer is correct

Concept

If X and Y are two mutually exclusive events, then P(X ∩ Y) = 0.

If X and Y are two events, then the conditional probability \({\bf{P}}\left( {{\bf{X}}/{\bf{Y}}} \right) = \frac{{{\bf{P}}\left( {{\bf{X}}; \cap ;{\bf{Y}}} \right)}}{{{\bf{P}}\left( {\bf{Y}} \right)}}\)

De Morgan’s Law

\(\overline {{\rm{A}} \cup {\rm{B}}} ; = {\rm{\bar A}} \cap {\rm{\bar B}}\)

\(\overline {{\rm{A}} \cap {\rm{B}}} = ;{\rm{\bar A}} \cup {\rm{\bar B}}\)

Calculation

Given A̅ and B̅ are mutually exclusive so P(A̅ ∩ B̅) = 0

Now,

P(A̅ ∪ B̅) = P(A̅) + P(B̅) - P(A̅ ∩ B̅)

\( \Rightarrow {\rm{P}}\left( {\overline {{\rm{A}} \cap {\rm{B}}} } \right) = 1 - {\rm{P}}\left( {\rm{A}} \right) + 1 - {\rm{P}}\left( {\rm{B}} \right)\)

⇒ 1 - P(A ∩ B) = 1  -P(A) + 1 - P(B)

⇒ P(A ∩ B) = P(A) + P(B) - 1

⇒ P(A ∩ B) = 0.1

The required conditional probability is \({\rm{P}}\left( {{\rm{A}}|{\rm{B}}} \right) = \frac{{{\rm{P}}\left( {{\rm{A;}} \cap {\rm{;B}}} \right)}}{{{\rm{P}}\left( {\rm{B}} \right)}} = \frac{1}{6}\)

Correct option is (2).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1436

National Defence Academy (NDA) 2019• Mathematics• Mathematics

Consider the following

Statements:

  1. The algebraic sum of deviations of a set of values from their arithmetic mean is always zero.
  2. Arithmetic mean > Median > Mode for a symmetric distribution.

Which of the above statements is/are correct?

  1. AOption (a)

    1 only

  2. BOption (b)

    2 only

  3. COption (c)

    Both 1 and 2

  4. DOption (d)

    Neither 1 nor 2

View answer & explanation

Correct answer

Answer: (a)A.

1 only

Why this answer is correct

Calculation

Suppose we have n observations X1, X2, ….., Xn.

The arithmetic mean \({\rm{\bar X}} = \frac{1}{{\rm{n}}}\mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\rm{X}}_{\rm{i}}}\)

Sum of deviations from arithmetic mean

\(= \mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} \left( {{{\rm{X}}_{\rm{i}}} - {\rm{\bar X}}} \right)\)

\( = \mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {{\rm{X}}_{\rm{i}}} - \mathop \sum \limits_{{\rm{i}} = 1}^{{\rm{i}} = {\rm{n}}} {\rm{\bar X}}\)

\(= {\rm{n\bar X}} - {\rm{n\bar X}}\)

Statement 1 is correct.

For a symmetric distribution mean, median and mode are equal i.e mean = median = mode. Hence, statement 2 is false.

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1437

National Defence Academy (NDA) 2019• Mathematics• Mathematics

Let the correlation coefficient between X and Y be 0.6. Random variables Z and W are defined as Z = X + 5 and \(W = \frac{Y}{3}\). What is the correlation coefficient between Z and W?

  1. AOption (a)

    0.1

  2. BOption (b)

    0.2

  3. COption (c)

    0.36

  4. DOption (d)

    0.6

View answer & explanation

Correct answer

Answer: (d)D.

0.6

Why this answer is correct

Concept

Let Z = aX + bY.

Expectation of Z = E(Z) = aE(X) + bE(Y)

Variance of Z = Var(Z) = a2Var(X) + b2Var(Y) + 2abρXYσXσY

ρXY = Correlation coefficient of X and Y.

\({{\rm{\rho }}_{{\rm{XY}}}} = \frac{{{\rm{Cov}}\left( {{\rm{X}},{\rm{;Y}}} \right)}}{{{{\rm{\sigma }}_{\rm{X}}}{{\rm{\sigma }}_{\rm{Y}}}}}\)

Covariance of X and Y = Cov(X, Y) = E(XY) - E(X)E(Y)

\({{\rm{\sigma }}_{\rm{X}}} = \sqrt {{\rm{Var}}\left( {\rm{X}} \right)} \)

\({{\rm{\sigma }}_{\rm{Y}}} = \sqrt {{\rm{Var}}\left( {\rm{Y}} \right)}\)

If Z = X + a where a is some constant, then Var(Z) = Var(X)

Covariance of X and Y = Cov(X, Y) = E(XY) - E(X)E(Y)

Calculation

Z = X + 5

⇒ Var(Z) = Var(X)

⇒ σZ = σX

\({\rm{W}} = \frac{{\rm{Y}}}{3}\)

\( \Rightarrow {\rm{Var}}\left( {\rm{W}} \right) = \frac{{{\rm{Var}}\left( {\rm{Y}} \right)}}{9}\)

\(\Rightarrow {\rm{;}}{{\rm{\sigma }}_{\rm{W}}} = \frac{{{{\rm{\sigma }}_{\rm{Y}}}}}{3}\)

Cov(Z, W)

= E(ZW) – E(Z)E(W)

\(= {\rm{E}}\left( {\frac{{{\rm{XY}}}}{3} + \frac{{5{\rm{Y}}}}{3}} \right) - {\rm{E}}\left( {{\rm{X}} + 5} \right){\rm{E}}\left( {\frac{{\rm{Y}}}{3}} \right)\)

\(= \frac{1}{3}{\rm{E}}\left( {{\rm{XY}}} \right) + \frac{5}{3}{\rm{E}}\left( {\rm{Y}} \right) - \frac{1}{3}{\rm{E}}\left( {\rm{X}} \right){\rm{E}}\left( {\rm{Y}} \right) - \frac{5}{3}{\rm{E}}\left( {\rm{Y}} \right)\)

\( = \frac{1}{3}\left( {{\rm{E}}\left( {{\rm{XY}}} \right) - {\rm{E}}\left( {\rm{X}} \right){\rm{E}}\left( {\rm{Y}} \right)} \right)\)

\(= {\rm{;}}\frac{1}{3}{\rm{Cov}}\left( {{\rm{X}},{\rm{;Y}}} \right)\)

Correlation coefficient of Z and W is

\({{\rm{\rho }}_{{\rm{ZW}}}} = \frac{{{\rm{Cov}}\left( {{\rm{Z}},{\rm{;W}}} \right)}}{{{{\rm{\sigma }}_{\rm{Z}}}{{\rm{\sigma }}_{\rm{W}}}}}\)

\(= \frac{{\frac{1}{3}{\rm{Cov}}\left( {{\rm{X}},{\rm{;Y}}} \right)}}{{{{\rm{\sigma }}_{\rm{X}}}\frac{{{{\rm{\sigma }}_{\rm{Y}}}}}{3}}}\)

\(= \frac{{{\rm{Cov}}\left( {{\rm{X}},{\rm{;;Y}}} \right)}}{{{{\rm{\sigma }}_{\rm{X}}}{{\rm{\sigma }}_{\rm{Y}}}}}\)

= ρXY

So ρZW = ρXY = 0.6

Correct option is (4).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1438

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If all the natural numbers between 1 and 20 are multiplied by 3, then what is the variance of the resulting series?

  1. AOption (a)

    99.75

  2. BOption (b)

    199.75

  3. COption (c)

    299.25

  4. DOption (d)

    399.25

View answer & explanation

Correct answer

Answer: (c)C.

299.25

Why this answer is correct

Concept

Sum of first n natural numbers is = \(\frac{{{\bf{n}}\left( {{\bf{n}} + 1} \right)}}{2}\)

Sum of square of first n natural numbers = \(\frac{{{\bf{n}}\left( {{\bf{n}} + 1} \right)\left( {2{\bf{n}} + 1} \right)}}{6}\)

Mean of first n natural numbers = \(\frac{1}{{\rm{n}}} \times {\rm{Sum;of;first;n;natural;numbers}} = \frac{{{\rm{n}} + 1}}{2}\)

If a constant k is multiplied with each of the first n natural numbers, then

\({\rm{Sum;is;}} = {\rm{k}}\frac{{{\rm{n}}\left( {{\rm{n}}; + ;1} \right)}}{2}\)

\({\rm{Mean;}} = {\rm{k}}\frac{{{\rm{n}}; + ;1}}{2}\)

The Variance is given as = \(\frac{1}{n}\left( {\mathop \sum \limits_{i = 1}^{i = n} {{\left( {i - {\rm{Mean}}} \right)}^2}} \right)\)

Calculation

Given that each number between 1 and 20 is multiplied by 3 so, the mean is

\(= 3\left( {\frac{{20; + ;1}}{2}} \right)\)

\(= \frac{{63}}{2}\)

The variance will be \(\frac{1}{{20}}\left( {\mathop \sum \limits_{i = 1}^{i = 20} {{\left( {3i - \frac{{63}}{2}} \right)}^2}} \right)\)

\(= \frac{1}{{20}}\mathop \sum \limits_{i = 1}^{i = 20} {\left( {3i - \frac{{63}}{2}} \right)^2}\)

\(= \frac{1}{{20}}\mathop \sum \limits_{i = 1}^{i = 20} \left( {9{i^2} + \frac{{3969}}{4} - 189i} \right)\)

\(= \frac{1}{{20}}\left[ {9\mathop \sum \limits_{i = 1}^{i = 20} {i^2} + \mathop \sum \limits_{i = 1}^{i = 20} \frac{{3969}}{4} - 189\mathop \sum \limits_{i = 1}^{i = 20} i} \right] \)

\(= \frac{1}{{20}}\left[ {9 \times \frac{{20 \times 21 \times 41}}{6} + 20 \times \frac{{3969}}{4} - 189 \times \frac{{20 \times 21}}{2}} \right]$ \)

\(= \frac{1}{{20}}\left( {25830 + 19845 - 39690} \right)\)

= 299.25

Correct option is (3).$

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1439

National Defence Academy (NDA) 2019• Mathematics• Mathematics

What is the probability that an interior point in a circle is closer to the centre than to the circumference?

  1. AOption (a)

    \(\frac{1}{4}\)

  2. BOption (b)

    \(\frac{1}{2}\)

  3. COption (c)

    \(\frac{3}{4}\)

  4. DOption (d)

    It cannot be determined

View answer & explanation

Correct answer

Answer: (a)A.

\(\frac{1}{4}\)

Why this answer is correct

Calculation

Consider two concentric circles as shown in the following figure:

Radius of the inner circle \(= {\rm{}}\frac{{\rm{r}}}{2}\)

Radius of the outer circle = r

Points lying inside the inner circle will definitely be closer to the centre than the circumference of the outer circle.

Possible region where these points lie = area of the inner circle = \(\frac{{{\rm{\pi }}{{\rm{r}}^2}}}{4}\)

Required probability \(= \frac{{{\rm{area;of;the;inner;circle}}}}{{{\rm{total;area}}}}\)

\(= \frac{{{\rm{\pi }}{{\rm{r}}^2}}}{{4; \times ;{\rm{\pi }}{{\rm{r}}^2}}}\)

\(= \frac{1}{4}\)

Correct option is (1).

NDA Mathematics 21 Apr 2019 Official Paper (Previous Year Paper)·Mathematics
View full paper

Question 1440

National Defence Academy (NDA) 2019• Mathematics• Mathematics

If A and B are two mutually exclusive events, then what is the probability of occurrence of either event A or event B?

  1. AOption (a)

    P(A) + P(B)

  2. BOption (b)

    P(A ∪ B')

  3. COption (c)

    P(A ∩ B)

  4. DOption (d)

    P(A) P(B)

View answer & explanation

Correct answer

Answer: (a)A.

P(A) + P(B)

Why this answer is correct

Concept

If A and B are two mutually exclusive events, then P(A ∩ B) = 0.

Calculation

Probability of occurrence of either event A or event B = P(A ∪ B)

Now,

P(A ∪ B) = P(A) + P(B) - P(A ∩ B)

⇒ P(A ∪ B) = P(A) + P(B)

Correct option is (1).

How should I prepare for Quantitative Aptitude in National Defence Academy (NDA)?

Solving previous year questions by topic helps identify recurring concepts, question patterns, and frequently asked sub-topics. Attempt each question, review the step-by-step explanation, and note down key formulas or facts for quick revision before the exam.